The perimeter of a rectangle is equal to twice the sum of its length and its width. one rectangle has a length of 13 1/8 inches and a perimeter of 42 1/2 inches. Write an equation that can be used to find the width, w, of the rectangle.
Answer by math_helper(2461)


Since perimeter p = 2w + 2l,
w = (p - 2l) / 2
Plugging in the given numbers:
w = (42.5 - 2*13.125) / 2 = 8.125 inches.